OpenStudy (anonymous):

allyj6 Group Title Write the equation of the line with a slope of 3/7 passing through the point (9,-2)

OpenStudy (texaschic101):

y = mx + b slope(m) = 3/7 (9,-2)...x = 9 and y = -2 now we sub -2 = 3/7(9) + b -2 = 27/7 + b -2 - 27/7 = b -14/7 - 27/7 = b -41 = b equation is : y = 3/7x - 41